OpenStudy (anonymous):

Find an explicit rule for the nth term of the sequence. 1, 3, 9, 27, ... (a) an = 1 • 3n + 1 (b) an = 3 • 1n - 1 (c) an = 1 • 3n - 1 (d) an = 1 • 3n

OpenStudy (anonymous):

\[1, 3, 9, 27, ...\] common ratio is 3 \[(3) \times 1 = 3\] \[(3) \times 3 = 9\] \[(3) \times 9 = 27\]

OpenStudy (anonymous):

\[a _{n} = a _{1} \times r ^{n -1}\] \[a _{n} = 1 \times 3 ^{n -1}\]

OpenStudy (anonymous):

so i believe you made a mistake and did not write your choices in the correct notation, but based on my intuition, I believe your answer is C.